on a map of the united states, the distance between Reno and Phoenix is 2.5 inches. the two cities are actually 768 miles apart. on this same map, what would be the distance between St.Louis and Jacksonville, two cities that are actually 919 miles apart?
Found 2 solutions by amoresroy, ewatrrr:
Answer by amoresroy(361)

on a map of the united states, the distance between Reno and Phoenix is 2.5 inches. the two cities are actually 768 miles apart. on this same map, what would be the distance between St.Louis and Jacksonville, two cities that are actually 919 miles apart?
Let x = the distance between St.Louis and Jacksonville on the map in inches
2.5/768 = x/919
x = 919/768*2.5
x = 2.99 inches

Answer by ewatrrr(24785)


Hi,
Let x represent the distance between St.Louis and Jacksonville on this same map
2.5in/768 = x/919
919*(2.5/768) = x
2.99in = 3in(to the nearest tenth of a inch) = x,
the distance between St.Louis and Jacksonville on this same map